# Realistic quantum network simulation for experimental BBM92 key distribution

### v1

@@ Abstract
Quantum key distribution (QKD) can provide secure key material between two parties without relying on assumptions about the computational power of an eavesdropper. QKD is performed over quantum links and quantum networks, systems which are resource-intensive to deploy and maintain. To evaluate and optimize performance prior to, during, and after deployment, realistic simulations with attention to physical realism are necessary. Quantum network simulators can simulate a variety of quantum and classical protocols and can assist in quantum network design and optimization by offering realism and flexibility beyond mathematical models which rely on simplifying assumptions and can be intractable to solve as network complexity increases. We use a versatile discrete event quantum network simulator to simulate the entanglement-based QKD protocol BBM92 and compare it to our experimental implementation and to existing theory. Furthermore, we simulate secure key rates in a repeater key distribution scenario for which no experimental implementations exist.
@@ I Introduction
Quantum key distribution (QKD) is a technique for creating a provably secure secret key between two parties by leveraging properties of quantum mechanics (such as monogamy of entanglement and the no-cloning theorem) bb84 ; QKDreviewGisin ; E91 ; decoystates ; twinfield . When used to create a one-time pad, QKD is information-theoretically secure, meaning it does not make any assumptions about the computational resources of an eavesdropper, and the maximum information the eavesdropper can possess can be determined. In contrast, commonly used classical public key cryptography systems like Rivest-Shamir-Adleman (RSA) RSA , elliptic curve cryptography (ECC) koblitz1987elliptic and even post-quantum cryptography (PQC) PQC offer only computational security. Their security is contingent upon the assumption that no efficient algorithms exist to invert the encryption using publicly shared information.
For RSA, an efficient algorithm to factor large numbers can break the encryption and render the encryption method insecure Shor1994 . An eavesdropper employing the harvest now decrypt later (HNDL) attack can intercept a message sent using RSA encryption and copy and store the encrypted message and public key until a future date when quantum computers have advanced enough to allow efficient decryption. Messages encrypted using PQC are similarly vulnerable. Although no efficient algorithms have yet been discovered to break the lattice-based cryptography underpinning PQC, HNDL attacks can still be levied on PQC systems in the same manner SIDHAttack .
@@
A variety of QKD protocols have been demonstrated over long distances yin2020 ; wang2022 ; Liu1000km with fiber multiplexing multiplex2018 ; Beutel2022 , via satellite ecker2021 ; satellite2025 and with secure authentication and integrated optical switches yongswitches2021 . When integrated in a quantum network with quantum memories, quantum error correction can be performed prior to measurements Feihureview2020 . However, QKD protocols vary in their security assumptions and vulnerability to attacks.
@@
Bennett-Brassard-Mermin 1992 (BBM92) is an entanglement-based QKD protocol in which Alice and Bob derive a shared secret key by measuring each entangled photon in a basis randomly selected from two orthogonal bases BBM92 . BBM92 lacks vulnerabilities present in other QKD schemes because it uses an entangled photon source to distribute the same maximally entangled Bell state to produce each key bit. In particular, QKD security assumptions including that the encoded state emitted from the source is basis-independent can be verified for the BBM92 protocol even if the source is controlled by an eavesdropper Feihureview2020 . High entanglement fidelity is a key requirement behind BBM92 security, and in deployed BBM92, maintaining a low quantum bit error rate (QBER) ensures the security of the protocol.
@@
Although analytical models exist and can provide useful estimates for secure key rates and other important metrics for BBM92 in simple quantum networks, the theory becomes intractable as the system complexity increases. Quantum network simulators can simulate parameter regimes which are not feasible to achieve with current-day devices, provide insights during experimental troubleshooting, and save the time and cost of a full physical experimental implementation. Particularly as quantum networks move towards full stack implementations, quantum network simulators can assist in simulation and optimization as quantum network protocols grow in complexity.
@@
In this paper, we accurately simulate BBM92 key exchange using the Aliro Quantum Network Simulator (AQNSim) alirosimulatorpub , a versatile and adaptable discrete event python simulation package for quantum networks with a high degree of customizability, physical accuracy, component modeling, and flexible abstraction. We compare our simulation results to our experimental implementation of BBM92 and to an existing numerical model for continuous-wave-pumped BBM92 NeumannPRA . We find that simulated results using AQNSim match our experiments as well as the theoretical model in regimes where the theoretical model is valid. Moreover, we leverage AQNSim to simulate entanglement generation rates and secure key distribution rates in a repeater network with multiple links, showing the utility of AQNSim for modeling scenarios where experiments may still be unfeasible and where analytical models may only exist under fixed assumptions.
@@ II BBM92 Experimental Setup
We implemented BBM92 key exchange using continuous-wave pumped, near-infrared polarization-entangled photons. In our setup, Alice and Bob generate a shared secret key by measuring photons emitted from an entangled photon source (see Figure 1). A narrowband entangled photon source emits polarization-entangled photons in the quantum state
@@
|$\ket{\Psi^{+}}=\frac{1}{\sqrt{2}}\left(\ket{HV}+\ket{VH}\right),$ (1)
where |$\ket{H}$ and |$\ket{V}$ form an orthonormal basis of horizontal and vertical single photon polarization states. The commercially-available source contains a laser which pumps two type-II spontaneous parametric down conversion (SPDC) quasi-phase-matched crystals in a Mach-Zehnder interferometer configuration (see Appendix A). The source emits photons at nominally 810 nm with a full-width half maximum (FWHM) of approximately 3 nm.
The entangled photons travel in single-mode fiber in a laboratory to each communication partner, Alice and Bob. At each communication partner, a fiber-based 50:50 beam splitter acts as a passive switch between an $HV$ basis measurement and an $AD$ basis measurement, where |$\ket{D}$$=\frac{1}{\sqrt{2}}\left(\ket{H}+\ket{V}\right)$ and |$\ket{A}=\frac{1}{\sqrt{2}}\left(\ket{H}-\ket{V}\right)$. Polarization compensation is performed to correct for polarization drift (see the measurement correlation matrix in Appendix B). At each communication partner, in-fiber polarization beam splitters route the received photon to one of four single-photon avalanche detectors for measurement. A time tagger with root-mean-square jitter of nominally 42 ps is used to collect and timetag detection events. Synchronization between Alice’s detectors and Bob’s detectors is necessary for BBM92 key exchange and is performed by characterizing the relative time delays for photons traveling to each detector (see Appendix A).
@@
To perform key exchange, time-tagged detection events are collected at each detector. Alice and Bob publicly exchange detection timestamps (adjusted for delays between detectors) labeled only by the randomly selected measurement basis. Alice and Bob then keep only the timestamped detection results which they both have measured in the same basis ("sifting"). A shared raw key between Alice and Bob is then distilled according to the detection results and anticipated correlations given the distributed Bell state (equation 1). The publicly exchanged basis information does not reveal any information about the secret key which would compromise the security, and any modification of the Bell state which is sent to Alice and Bob will result in a detectable increase in QBER. In this work, the QBER is calculated by comparing Alice and Bob’s entire raw keys. In a practical deployment, a fraction of the key is sacrificed to estimate the QBER of the entire key. Following the estimation of the QBER, error correction can be performed (using, for instance, low density parity check codes LDPC ) followed by error verification and authentication. As a final step, privacy amplification can minimize the information an eavesdropper might extract Feihureview2020 .
III Discrete Event Simulation of Quantum Networks
@@
Discrete event simulation of quantum networks combines a full quantum mechanical treatment of qubits with event-based logic and timing for quantum and classical signals. Discrete event quantum network simulators such as SeQUeNCe sequencepaper , QuISP quisp , NetSquid netsquid , and others QNSimulatorsReview have accelerated quantum network protocol design. For existing discrete event quantum network simulators, quantum memories, qubits, and quantum gates are often highly abstracted from their physical representations or else specific to a given physical implementation. In contrast, AQNSim is a quantum network discrete event simulation python package which provides realistic modeling of quantum optics, from Fock states to beam splitters. Depending on the experiment and application, this realism can be critical to accurately modeling qubit behavior in a quantum network. For BBM92, hardware specifications like timing jitter, detector dead times, and dark counts all are necessary for accurate modeling of key rates and QBERs and can be incorporated using AQNSim.
@@
AQNSim can simulate quantum key distribution protocols, entanglement swapping and purification, quantum gates and quantum operators, quantum optical components including beam splitters and waveplates, nonlinear optical processes such as spontaneous down conversion, and more at a user-specified level of abstraction. The event-based nature of AQNSim supports the simulation of classical messages and communication on the network. When needed for increased performance, AQNSim also has a multi-shot capability, allowing many measurements to be performed in parallel by calculating the pre-measurement quantum state once and sampling from it many times.
@@
|$\ket{W(F)}=\frac{4F-1}{3}\ket{\Psi^{+}}\bra{\Psi^{+}}+\frac{1-F}{3}I,$ (2)
where $F$ is the fidelity between the quantum state measured at Alice and Bob and the desired Bell state ($F=|\innerproduct{\Psi^{+}}{W(F)}|^{2}$). When the probability of measuring an accidental coincidence between photons from different entangled pairs is small compared to the probability of measuring a coincidence between photons from the same entangled pair, $F$ is related to the QBER of the measurement results in each basis ($X$, $Y$, $Z$) by fidelityqber
@@
For the simulations, we find $F$ from the measured QBER from our experiments.
@@
IV Theoretical Model
Existing theoretical models can predict raw key rates, secure key rates, and QBERs for continuous-wave-pumped BBM92 NeumannPRA . In this section, we follow these models to produce theoretical estimates of key rates and QBERs which will then be used to verify our BBM92 experiment and simulation.
@@ IV Theoretical Model
Photon coincidence rates between Alice and Bob are predicted by considering a theoretical histogram showing differences between arrival times between detections at Alice and detections at Bob. The total number of coincidences in a coincidence window $t_{c}$ is thus given by the integral of the coincidence histogram from $\mu-t_{c}/2$ to $\mu+t_{c}/2$, where $\mu$ is the offset of the histogram peak along the x-axis and $t_{c}$ is the coincidence window size. Coincidence rates can be measured as this integral divided by total acquisition time and $t_{c}$.
@@
The probability of the entangled photon source emitting $n$ photon pairs in a given coincidence window follows a Poisson distribution QKDreviewGisin . For a Poisson point process, the probability of $n$ events occurring in a time period given the average rate of event per time period $\lambda$ is
@@
Figure 3: (a) Raw key rates (with relative errors as inset plot) and (b) QBERs (with relative errors as inset plot) for theory, experiment, and AQNSim simulation of BBM92 key exchange. Relevant experimental parameters used in simulations are provided in Table 1. The AQNSim key rates and QBERs are obtained from a randomly seeded simulation, re-binned at each coincidence window, averaged over 4,000,000 shots. The experimental key rates and QBERs are obtained from a single experimental run, with the data re-binned at each coincidence window size.
@@
To find the probability of error per bit, or QBER, several different potential sources of error must be considered. These sources of error include the error rate due to imperfections in the optical setup ($R_{\text{opt, err}}$), and the error rate from accidental coincidences that do not produce valid key bits ($R_{\text{acc, err}}$).
Imperfections in the optical setup contribute to error as in experimental reality, the quantum state which reaches Alice and Bob’s photon detectors will not necessarily have perfect fidelity with the desired quantum state from equation 1. Defects in measurement optics like waveplates or polarization beam splitters, imperfect polarization drift compensation, fluctuations in temperature, and small misalignments in quasi-phase matching within the integrated source will lower the visibility of the measured quantum state.
@@
Error correction protocols in deployed BBM92 systems will consume valid key bits in order to detect and reconcile key errors between Alice and Bob. For an equal probability of measurement in either the $HV$ or $AD$ basis, and for a conservative assumption on the amount of key consumed to detect and rectify errors which accounts for finite key block sizes in the error correction process, we can predict the asymptotic secure key rate as Elkouss2011
@@
The asymptotic secure key rate is the maximum key rate in the limit of infinitely long keys. For finite key lengths, the value of the QBER (needed for choosing the most efficient error correction code and for privacy amplification) must be estimated by sampling bits from the key. From the estimated QBER the secure key rate for a finite key can then be probabilistically bounded.
V BBM92 Simulation, Experiment, and Theory Comparison
@@
Figure 4: Asymptotic BBM92 secure key rate for simulation, theory, and experiment for swept parameters (a) loss per link, (b) detector dead time, (c) detection resolution, (d) dark count rate per communication partner, (e) source brightness, and (f) visibility. The coincidence window is 2 ns and all other parameters are given in Table 1.
Raw key rates and QBERs for the experiment for different coincidence window durations are plotted in Figure 3. The experimental data plotted in red is re-binned at each coincidence window to calculate the raw key rate and QBER. Using methods described in Appendix C, parameters such as internal source brightness, loss per link, detection resolution, and visibility are estimated from the experimental data. Other relevant parameters, including detector dead time, are obtained from equipment data sheets. These parameters are summarized in Table 1.
@@
Theory, experiment, and simulation show excellent agreement. Relative errors for the raw key rate and QBER for both the theoretical model as well as for AQNSim simulation are plotted in insets in Figure 3 for coincidence windows of 100 picoseconds or larger. At smaller coincidence windows the error bars are larger because the number of bits in the simulated and experimental key decreases. AQNSim simulation predicts the QBER within 1 $\sigma$ error margins and predicts the raw key rates within a few percent. The deviation of simulation from experiment is near the tolerance expected given the uncertainties on the estimated parameters used in simulation. The inset in Figure 3a shows the theoretical model displaying a growing deviation for the predicted raw key rate from the experimentally measured raw key rate as coincidence windows increase in size (beginning around coincidence windows greater than 1 microsecond). Deviation of experimental results from theory occurs from larger coincidence windows leading to more accidental coincidences which eliminate true coincidences by falling in the same coincidence window.
AQNSim can provide helpful design verification for BBM92 asymptotic secure key rates for a wide range of experiment parameter regimes. We simulated the asymptotic secure key rate for our BBM92 setup using AQNSim and plotted the rate against different varying parameters including source brightness, dark count rate, detector dead time, detection resolution, loss per link, and visibility (see Figure 4). The secure key rate is calculated from the raw key rate, QBER, and equation 18. In Figure 4 the AQNSim-simulated secure key rates match theoretical predictions and experiment. AQNSim can accurately predict the secure key rate for BBM92 and can assist in finding trends and tolerances for experimental parameters.
@@
VI BBM92 simulation and theory via repeaters
Photons traveling in a fiber experience loss, and the probability of photon loss increases exponentially with the length of the fiber. This photon loss together with detector dark counts limit the maximum distance over which quantum secure communication can be deployed. To counteract this loss, trusted relay nodes have been proposed for prepare-and-measure QKD protocols like BB84 elliott2005 ; trustedRelayNodePRA . Trusted relay nodes can extend the distance over which prepare-and-measure QKD protocols can be performed, but their security relies on the assumption that an eavesdropper cannot access and manipulate the relay node.
To enable quantum secure communication between two parties over long distances without security assumptions that the eavesdropper cannot access relay nodes, quantum repeaters have been proposed Briegel1998 ; repeaterReview ; atomicensembles . In a quantum repeater network, the quantum channel between Alice and Bob is divided into segments (elementary links), with a quantum repeater node placed between each elementary link. Entanglement is first generated across the adjacent elementary links and stored in quantum memories, and then established between Alice and Bob by applying entanglement swapping to each quantum memory in each quantum repeater node. To obtain the shared secret bit, Alice and Bob then each measure the qubit stored in their quantum memory.
Figure 5: (a) Entanglement creation using a repeater containing two end nodes (Alice and Bob), one central repeater node with two stored qubits capable of two-qubit gate operations ($r=1)$, and two Bell state measurement stations (BSM). (b) The average entanglement generation rate (in units of $c/L_{0}$) for different elementary link loss values from simulation (using AQNSim with 1000 shots) and theory for a two-link repeater.
We simulate the entanglement generation rates and secure key rates for BBM92 in a scenario where entanglement is distributed between two end nodes via a repeater network containing $r$ repeater nodes. The repeater protocol begins after a synchronization message is sent by Alice to each node indicating the timestamp when all protocols should be started. Each node then independently runs its own protocol to first generate entanglement across the elementary links. For each elementary link, the associated nodes each entangle a photonic qubit with a stationary qubit in their quantum memory, and then send the photonic qubit to a Bell state measurement (BSM) station located in the center of the elementary link. The photonic qubits are measured at the Bell state measurement station and the results transmitted to the adjacent nodes.

### v3

@@ Abstract
Quantum key distribution (QKD) can provide secure key material between two parties without relying on assumptions about the computational power of an eavesdropper. QKD is performed over quantum links and quantum networks, systems which are resource-intensive to deploy and maintain. To evaluate and optimize performance prior to, during, and after deployment, accurate simulations with attention to physical realism are necessary. Quantum network simulators can simulate a variety of quantum and classical protocols and can assist in quantum network design and optimization by offering realism and flexibility beyond mathematical models which rely on simplifying assumptions and can be intractable to solve as network complexity increases. We use a versatile discrete event quantum network simulator to simulate the entanglement-based QKD protocol BBM92 and compare it to our experimental implementation and to existing theory. We find the discrete event quantum network simulator can match experimental key rates and error rates with a lower mean squared error than analytical theory. Furthermore, we simulate secure key rates in a repeater key distribution scenario for which no experimental implementations exist and find agreement between simulation and analytical theory. Hence, we demonstrate discrete event simulators can meet needs in quantum network simulations which cannot be filled solely by experiment or theory: discrete event simulators can accurately simulate QKD protocols and match experiments in regimes where theoretical models may require more simplifying assumptions, and they can match theoretical models in the opposite scenario where experiments have not yet been performed but theoretical models exist.
@@ I Introduction
Quantum key distribution (QKD) is a technique for creating a provably secure secret key between two parties by leveraging properties of quantum mechanics (such as monogamy of entanglement and the no-cloning theorem) [6, 17, 14, 21, 22]. When used to create a one-time pad, QKD is information-theoretically secure, meaning it does not make any assumptions about the computational resources of an eavesdropper, and the maximum information the eavesdropper can possess can be determined. In contrast, commonly used classical public key cryptography systems like Rivest-Shamir-Adleman (RSA) [28], elliptic curve cryptography (ECC) [18] and even post-quantum cryptography (PQC) [3] offer only computational security. Their security is contingent upon the assumption that no efficient algorithms exist to invert the encryption using publicly shared information.
For RSA, an efficient algorithm to factor large numbers can break the encryption and render the encryption method insecure [33]. An eavesdropper employing the harvest now decrypt later (HNDL) attack can intercept a message sent using RSA encryption and copy and store the encrypted message and public key until a future date when quantum computers have advanced enough to allow efficient decryption. Messages encrypted using PQC are similarly vulnerable. Although no efficient algorithms have yet been discovered to break the lattice-based cryptography underpinning PQC, HNDL attacks can still be levied on PQC systems in the same manner [10].
@@
A variety of QKD protocols have been demonstrated over long distances [35, 41, 36, 20] with fiber multiplexing [37, 8], via satellite [13, 19] and with secure authentication and integrated optical switches [40]. When integrated in a quantum network with quantum memories, quantum error correction can be performed prior to measurements [39]. However, QKD protocols vary in their security assumptions and vulnerability to attacks.
@@
Bennett-Brassard-Mermin 1992 (BBM92) is an entanglement-based QKD protocol whereby Alice and Bob derive each bit in a shared secret key by measuring each photon from an entangled Bell pair in a basis randomly selected from two orthogonal bases [5]. BBM92 lacks vulnerabilities present in other QKD schemes because it uses an entangled photon source to distribute the same maximally entangled Bell state to produce each key bit. In particular, QKD security assumptions including that the encoded state emitted from the source is basis-independent can be verified for the BBM92 protocol even if the source is controlled by an eavesdropper [39]. High entanglement fidelity is a key requirement behind BBM92 security, and in deployed BBM92, maintaining a low quantum bit error rate (QBER) ensures the security of the protocol.
@@
Although analytical models exist and can provide useful estimates for secure key rates and other important metrics for BBM92 in simple quantum networks, the theory becomes intractable as the system complexity increases. In contrast, quantum network simulators can simulate parameter regimes which are not feasible to achieve with current-day devices, provide insights during experimental troubleshooting, and save the time and cost of a full physical experimental implementation. Particularly as quantum networks move towards full stack implementations, quantum network simulators can assist in simulation and optimization as quantum network protocols grow in complexity.
@@
In this paper, we accurately simulate BBM92 key exchange using the Aliro Quantum Network Simulator (AQNSim) [1], a versatile and adaptable discrete event Python simulation package for quantum networks. We choose to use AQNSim for our simulations because of its high degree of customizability, physical accuracy, detailed component modeling, and flexible abstraction. While we do not perform a systematic comparison across all available simulators, we use AQNSim as a reference implementation to show that a Python-based discrete event quantum network simulation framework can be used to accurately and realistically simulate quantum network protocols. We compare our simulation results to our experimental implementation of BBM92 and to an existing numerical model for continuous-wave-pumped BBM92 [25]. We find that simulated results using AQNSim match our experiments as well as the theoretical model in regimes where the theoretical model is valid. In regimes where assumptions required for the theoretical model fail, our simulations match experimental data with a lower average mean squared error. Moreover, we leverage AQNSim to simulate entanglement generation rates and secure key distribution rates in a repeater network with multiple links, showing the utility of AQNSim for modeling scenarios where experiments may still be unfeasible and where analytical models may only exist under fixed assumptions.
@@ II BBM92 Experimental Setup
We implemented BBM92 key exchange using continuous-wave pumped, near-infrared polarization-entangled photons. In our experimental setup, Alice and Bob generate a shared secret key by measuring photons emitted from an entangled photon source (see Figure 1). A narrowband entangled photon source emits polarization-entangled photons in the quantum state
@@
[…]
@@
To perform key exchange, time-tagged detection events are collected at each detector. Alice and Bob publicly exchange detection timestamps (adjusted for delays between detectors) labeled only by the randomly selected measurement basis. Alice and Bob then keep only the timestamped detection results which they both have measured in the same basis ("sifting"). A shared raw key between Alice and Bob is then distilled according to the detection results and anticipated correlations given the distributed Bell state (equation 1). The publicly exchanged basis information does not reveal any information about the secret key which would compromise the security, and any modification of the Bell state which is sent to Alice and Bob will result in a detectable increase in QBER. In this work, the QBER is calculated by comparing Alice and Bob’s entire raw keys. In a production-ready deployment, a fraction of the key is sacrificed to estimate the QBER of the entire key. Following the estimation of the QBER, error correction can be performed (using, for instance, low density parity check codes [24]) followed by error verification and authentication. As a final step, privacy amplification can minimize the information extractable by an eavesdropper [39].
III Discrete Event Simulation for Quantum Networks
@@
Discrete event simulation of quantum networks combines a full quantum mechanical treatment of qubits with event-based logic and timing for quantum and classical signals. Discrete event quantum network simulators such as SeQUeNCe [38], QuISP [30], NetSquid [11], and others [4] have accelerated quantum network protocol design. For existing discrete event quantum network simulators, quantum memories, qubits, and quantum gates are often highly abstracted from their physical representations or else specific to a given physical implementation. In comparison, AQNSim is a quantum network discrete event simulation Python package which provides realistic modeling of quantum optics with customizable abstraction. Depending on the experiment and application, this realism can be critical to accurately modeling qubit behavior in a quantum network. For BBM92, hardware specifications like timing jitter, detector dead times, and dark counts all are necessary for accurate modeling of key rates and QBERs and can be incorporated using AQNSim.
@@
AQNSim can simulate the physics and timing of quantum key distribution protocols, entanglement swapping, quantum gates and quantum operators, quantum optical components including beam splitters and waveplates, and more at a user-specified level of abstraction. The event-based nature of AQNSim supports the simulation of classical messages and communication on the network. When needed for increased performance, AQNSim also has a multi-shot capability, allowing many measurements to be performed in parallel by calculating the pre-measurement quantum state once and sampling from it many times.
@@
[…]
where $F$ is the fidelity between the quantum state measured at Alice and Bob and the desired Bell state.
We adopt the Werner state model because we assume that our generated entangled photon pairs are approximately subject to isotropic depolarizing noise. Isotropic depolarizing noise replaces the maximally entangled state with the maximally mixed state with some probability and is often used to approximate generalized phenomenological qubit noise without requiring a full determination of the exact noise channel that the qubit experiences [26]. We find this approximation reproduces the experimental data with a high degree of accuracy without requiring full process tomography.
When the probability of measuring an accidental coincidence between photons from different entangled pairs is small compared to the probability of measuring a coincidence between photons from the same entangled pair, $F$ is related to the QBER of the measurement results in each basis ($X$, $Y$, $Z$) by [12]
@@
For the simulations, we find $F$ from the measured QBER from our experiments (see Appendix C for further details).
@@
IV Theoretical Model for BBM92 Key Generation
Existing theoretical models can predict raw key rates, secure key rates, and QBERs for continuous-wave-pumped BBM92 [25]. The theoretical model for key generation rates for continuous-wave-pumped BBM92 is a modification of the theoretical model for key generation rates for pulsed-pump BBM92 using parametric down conversion sources [23]. In this section, we follow these models to produce theoretical estimates of key rates and QBERs which will then be compared with our BBM92 experimental results and simulation results.
@@ IV Theoretical Model for BBM92 Key Generation
Photon coincidence rates between Alice and Bob are predicted in the analytical model by considering a theoretical histogram showing differences between arrival times between detections at Alice and detections at Bob. The total number of coincidences in a coincidence window $t_{c}$ is thus given by the integral of the coincidence histogram from $\mu-t_{c}/2$ to $\mu+t_{c}/2$, where $\mu$ is the offset of the histogram peak along the x-axis and $t_{c}$ is the coincidence window size. Coincidence rates can be measured as this integral divided by total acquisition time and $t_{c}$.
@@
The probability of the entangled photon source emitting $n$ photon pairs in a given coincidence window follows a Poisson distribution [17]. For a Poisson point process, the probability of $n$ events occurring in a time period given the average rate of event per time period $\lambda$ is
@@
To find the probability of error per bit, or QBER, several different potential sources of error must be considered. These sources of error include the error due to imperfections in the optical setup ($R_{\text{opt, err}}$), and the error from accidental coincidences that do not produce valid key bits ($R_{\text{acc, err}}$).
Imperfections in the optical setup contribute to error by causing the quantum state which reaches Alice and Bob’s photon detectors to deviate from the desired quantum state (equation 1). Defects in measurement optics like waveplates or polarization beam splitters, imperfect polarization drift compensation, fluctuations in temperature, and small misalignments in quasi-phase matching within the integrated source will lower the visibility of the measured quantum state.
@@
Error correction protocols in deployed BBM92 systems will consume valid key bits in order to detect and reconcile key errors between Alice and Bob. With error correction, secure key distribution can still be performed in the presence of errors and a theoretical bound on the maximum secure key rate given an infinitely long key length, the asymptotic secure key rate, can be determined [32]. For an equal probability of measurement in either the $HV$ or $AD$ basis, and for a conservative assumption on the amount of key consumed to detect and rectify errors which accounts for finite key block sizes in the error correction process, we can predict the asymptotic secure key rate as [15]
@@
For finite key lengths, the value of the QBER (needed for choosing the most efficient error correction code and for privacy amplification) must be estimated by sampling bits from the key. From the estimated QBER the secure key rate for a finite key can then be probabilistically bounded.
V Comparing Simulation, Experiment, and Theory for BBM92 Key Generation
Figure 3: (a) Raw key rates (with relative errors as inset plot) and (b) QBERs (with relative errors as inset plot) for theory, experiment, and AQNSim simulation of BBM92 key exchange. Relevant experimental parameters used in simulations are provided in Table 1. The AQNSim key rates and QBERs are obtained from a randomly seeded simulation, re-binned at each coincidence window, averaged over 4,000,000 shots. The experimental key rates and QBERs are obtained from a single experimental run, with the data re-binned at each coincidence window size. Root mean squared errors are displayed for AQNSim and experiment and for theory and experiment for coincidence windows greater than 100 picoseconds.
@@ V Comparing Simulation, Experiment, and Theory for BBM92 Key Generation
Figure 4: Asymptotic BBM92 secure key rate for simulation, theory, and experiment for swept parameters (a) loss per link, (b) detector dead time, (c) detection resolution, (d) dark count rate per communication partner, (e) source brightness, and (f) visibility. The coincidence window for all plots is 2 nanoseconds and all other parameters are given in Table 1.
Raw key rates and QBERs for the experiment for different coincidence window durations are plotted in Figure 3. The experimental data plotted in red is re-binned at each coincidence window to calculate the raw key rate and QBER. For the AQNSim model and for the analytical model of the BBM92 key exchange, parameters such as internal source brightness, loss per link, detection resolution, and entangled photon source and optical setup visibility are estimated from the experimental data using methods described in Appendix C. The experimentally measured QBER at 1 nanosecond is used to set the fidelity of the Werner state for the simulated and analytical models and this fidelity is verified with independent entangled photon source visibility measurements. Other relevant parameters, including detector dead time, are obtained from equipment data sheets. These parameters are summarized in Table 1.
@@
Theory, experiment, and simulation show excellent agreement. Relative errors for the raw key rate and QBER for both the theoretical model as well as for AQNSim simulation are plotted in insets in Figure 3 for coincidence windows of 100 picoseconds or larger. At smaller coincidence windows the error bars are larger because the number of bits in the simulated key and in the experimentally derived key decreases. AQNSim simulation predicts the QBER within 1 $\sigma$ error margins and predicts the raw key rates within a few percent. The deviation of simulation from experiment is near the tolerance expected given the uncertainties on the estimated parameters used in simulation. The inset in Figure 3a shows the theoretical model displaying a growing deviation for the predicted raw key rate from the experimentally measured raw key rate as coincidence windows increase in size (beginning around coincidence windows greater than 1 microsecond). Deviation of experimental results from theory occurs from larger coincidence windows leading to more accidental coincidences which eliminate true coincidences by falling in the same coincidence window as well as from an overestimation in accidental coincidences in the theoretical model by counting true coincidences as accidental.
AQNSim can predict BBM92 asymptotic secure key rates for a wide range of equipment specifications and operating regimes. We simulate the asymptotic secure key rate for our BBM92 setup using AQNSim and plot the rate against different experimental parameters including source brightness, dark count rate, detector dead time, detection resolution, loss per link, and visibility (see Figure 4). Here, the secure key rate is calculated from the raw key rate, QBER, and equation 18. In Figure 4 the AQNSim-simulated secure key rates match theoretical predictions and experiment. AQNSim can accurately predict the secure key rate for BBM92 and can assist in finding trends and tolerances for experimental parameters.
@@
VI Comparing Simulation and Theory for BBM92 in Quantum Repeaters
Photons traveling in a fiber experience loss, and the probability of photon loss increases exponentially with the length of the fiber. This photon loss together with detector dark counts limit the maximum distance over which quantum secure communication can be deployed. To counteract this loss, trusted relay nodes have been proposed for prepare-and-measure QKD protocols like BB84 [16, 34]. Trusted relay nodes can extend the distance over which prepare-and-measure QKD protocols can be performed, but their security relies on the assumption that an eavesdropper cannot access and manipulate the relay node.
To enable quantum secure communication between two parties over long distances without security assumptions requiring trusted relay nodes, quantum repeaters have been proposed [9, 2, 29]. In a quantum repeater network, the quantum channel between Alice and Bob is divided into segments (elementary links), with a quantum repeater node placed between each elementary link. Entanglement is first generated across the adjacent elementary links and stored in quantum memories, and then established between Alice and Bob by applying entanglement swapping to each quantum memory in each quantum repeater node. To obtain the shared secret bit, Alice and Bob then each measure the qubit stored in their quantum memory.
Figure 5: (a) Entanglement creation using a repeater containing two end nodes (Alice and Bob), one central repeater node with two stored qubits capable of two-qubit gate operations ($r=1)$, and two Bell state measurement stations (BSM). (b) The average entanglement generation rate (in units of $c/L_{0}$) for different elementary link loss values calculated from simulation (using AQNSim with 1000 shots) and theory for a two-link repeater.
We simulate the entanglement generation rates and secure key rates for BBM92 in a scenario where entanglement is distributed between two end nodes via a repeater network containing $r$ repeater nodes. The repeater protocol begins after a synchronization message is sent by Alice to each node indicating the timestamp when all protocols should be started. Each node then independently runs its own protocol to first generate entanglement across the elementary links. For each elementary link, the associated nodes each entangle a photonic qubit with a stationary qubit in their quantum memory, and then send the photonic qubit to a Bell state measurement station (BSM) located in the center of the elementary link. The photonic qubits are measured at the Bell state measurement station and the results transmitted to the adjacent nodes.